OpenStudy (anonymous):

the population of a country is currently at 56 million and is forecast to rise by 3.7 % each year. it is capable of producing 2500 million units of food each year, and it is estimated that each member of the population requires a minimum of 65 unit of food each year. at the moment , the extra food needed to satisfy this requirement is imported , but the government decides to increase food production at a constant rate each year , with the aim of making the country self sufficient after 10 years. find the annual rate of growth requied to achieve this.

OpenStudy (anonymous):

First, write the equation for population growth.

OpenStudy (anonymous):

56(1+3.7/100) ?

OpenStudy (anonymous):

It should be (population)1.037

OpenStudy (anonymous):

37 % the population right ?

OpenStudy (anonymous):

It's the same as what you said, but accounts for the varying pop. It's 103.7% --> 1.037

OpenStudy (anonymous):

Now, find what the population will be in 10 years.

OpenStudy (anonymous):

what about food?

OpenStudy (anonymous):

56(1+3.7/100)^10

OpenStudy (anonymous):

Since we want to find out how much food we'll need in 10 years, we need to know how many people there will be in 10 years.

OpenStudy (anonymous):

56(1+3.7/100)^10=86.60 right>

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

for 56(1+3.7/100)=58

OpenStudy (anonymous):

I got 80.5

OpenStudy (anonymous):

yea you have right !

OpenStudy (anonymous):

Okay, so we have 80.5 million people in 10 years. How much food will they need to eat?

OpenStudy (anonymous):

minimum required of 65 unit for food each year and it is capaple to producing 2500 million unit of food

OpenStudy (anonymous):

65 unit for food each year per person right? So what will we do to find the total amount?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

the anwer we find divide by the food per person?

OpenStudy (anonymous):

Actually, we will multiply so there will be 65 units for each of the millions of people.

OpenStudy (anonymous):

80.5 mutiplie by 65?

OpenStudy (anonymous):

Remember It's not just 80.5, it's 80.5 million.

OpenStudy (anonymous):

yea i know but the anwser will in million but is correct that i shoudl 80.5 multiplie by 65=5232.5 million ?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

we need annual rate of growth required

OpenStudy (anonymous):

We know how much we need, so now we need a function that will get us from the current annual growth rate to 5232.5 million food units.

OpenStudy (anonymous):

so 5232.5( 1+r/100)?

OpenStudy (anonymous):

Not exactly. 5232.5 is the result we already have. It's where we need to get to. It will be to the right of the "=".

OpenStudy (anonymous):

What's the current annual food production?

OpenStudy (anonymous):

each year annual food

OpenStudy (anonymous):

yes. What is the food production currently?

OpenStudy (anonymous):

they can produce 2500 million unit of food

OpenStudy (anonymous):

Right, now what will get that equal to 5232.5 in ten years?

OpenStudy (anonymous):

here i dont know ^^

OpenStudy (anonymous):

It's the compound interest formula. What's the beginning amount you have?

OpenStudy (anonymous):

56million rise 3.7 % i know it is compound interest formula but how i can do for food if i dont have rate

OpenStudy (anonymous):

ah maybe because food per populatioN

OpenStudy (anonymous):

EXACTLY! You don't have the food rate. When you don't know what something is, we call it a variable x

OpenStudy (anonymous):

so 2500(1+r/100) then we can do for 10 year?

OpenStudy (anonymous):

Right. What do you do with the ten? Don't forget to set it equal to 5232.5.

OpenStudy (anonymous):

ahh kk so 2500(1+r/100)^10=5232.5

OpenStudy (anonymous):

Yep, that's the equation!

OpenStudy (anonymous):

You just need to solve for r.

OpenStudy (anonymous):

so then 5232.5 /2500=(1+r/100)^10 ln(5232.5/2500)=10ln(1+r/100) ln(5232.5/2500)/10=ln(1+r/100_ i do expenential? for both

OpenStudy (anonymous):

then x anwser-1multiplied 100= r?

OpenStudy (anonymous):

Get rid of the 2500 first. Then take log 10 of both sides. What does that give you?

OpenStudy (anonymous):

i found the final anwser 7.66 %

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

THANK YOU :D awnsome how i can give you medal? i am new here ^^

OpenStudy (anonymous):

You click "Best Response". No problem.

OpenStudy (anonymous):

It should be right next to the medal on one of my responses.